# Capturing a Moving Target by Two Robots in the F2F Model

### v1

@@ 5.1. Target Moving Away from the Origin
$8M^{7}+64(\log\log M+3)\cdot M^{12}\cdot\log^{2}M\cdot\left(\frac{1-v}{d}+\frac{2}{d}\right),$
@@
$\sum_{j=0}^{i-1}x_{j}$ $=\frac{x_{i}(u_{i}-v_{i})-2^{g_{i}}}{v_{i}}\implies\sum_{j=0}^{i}x_{j}=\frac{x_{i+1}(u_{i+1}-v_{i+1})-2^{g_{i+1}}}{v_{i+1}}.$
@@
Thus, based on Equation (6), we have the following:
@@
$CR$ $=\frac{\sum_{j=0}^{i-1}x_{j}+\frac{d_{i}+v\sum_{j=0}^{i-1}x_{j}}{u_{i}-v}+\frac{2\sum_{j=0}^{i-1}u_{j}x_{j}+2u_{i}\left(\frac{d_{i}+v\sum_{j=0}^{i-1}x_{j}}{u_{i}-v}\right)}{1-u_{i}}}{\frac{d}{1-v}}$
$+\frac{\frac{2\sum_{j=0}^{i-1}u_{j}x_{j}+2u_{i}\left(\frac{d_{i}+v\sum_{j=0}^{i-1}x_{j}}{u_{i}-v}\right)+2v\sum_{j=0}^{i-1}u_{j}x_{j}+2vu_{i}\left(\frac{d_{i}+v\sum_{j=0}^{i-1}x_{j}}{u_{i}-v}\right)}{(1-u_{i})(1-v)}}{\frac{d}{1-v}}$
$=\frac{\frac{u_{i}\sum_{j=0}^{i-1}x_{j}+d_{i}}{u_{i}-v}+\frac{2u_{i}\sum_{j=0}^{i-1}u_{j}x_{j}-2v\sum_{j=0}^{i-1}u_{j}x_{j}+2u_{i}d_{i}+2vu_{i}\sum_{j=0}^{i-1}x_{j}}{(1-u_{i})(1-v)(u_{i}-v)}}{\frac{d}{1-v}}$
$+\frac{\frac{2u_{i}\sum_{j=0}^{i-1}u_{j}x_{j}-2v\sum_{j=0}^{i-1}u_{j}x_{j}+2u_{i}d_{i}+2u_{i}v\sum_{j=0}^{i-1}x_{j}}{(1-u_{i})(1-v)(u_{i}-v)}}{\frac{d}{1-v}}$
$\leq\frac{u_{i}\sum_{j=0}^{i-1}x_{j}+d_{i}-u_{i}v\sum_{j=0}^{i-1}x_{j}-vd_{i}-(u_{i})^{2}\sum_{j=0}^{i-1}x_{j}-u_{i}d_{i}+u_{i}^{2}v\sum_{j=0}^{i-1}x_{j}}{d_{i}(1-u_{i})(u_{i}-v)}$
$+\frac{d_{i}u_{i}v+2u_{i}\sum_{j=0}^{i-1}u_{j}x_{j}-2v\sum_{j=0}^{i-1}u_{j}x_{j}+2u_{i}d_{i}+2vu_{i}\sum_{j=0}^{i-1}x_{j}}{d(1-u_{i})(u_{i}-v)}$
$+\frac{2u_{i}\sum_{j=0}^{i-1}u_{j}x_{j}-2v\sum_{j=0}^{i-1}u_{j}x_{j}+2u_{i}d_{i}+2u_{i}v\sum_{j=0}^{i-1}x_{j}}{d(1-u_{i})(u_{i}-v)}$
$\leq\frac{u_{i}\sum_{j=0}^{i-1}x_{j}+d_{i}-d_{i}v+3u_{i}d_{i}-u_{i}^{2}\sum_{j=0}^{i-1}x_{j}+u_{i}^{2}v\sum_{j=0}^{i-1}x_{j}+d_{i}u_{i}v}{d(1-u_{i})(u_{i}-v)}$
$+\frac{4u_{i}\sum_{j=0}^{i-1}u_{j}x_{j}-4v\sum_{j=0}^{i-1}u_{j}x_{j}-u_{i}v\sum_{j=0}^{i-1}x_{j}+4u_{i}v\sum_{j=0}^{i-1}x_{j}}{d(1-u_{i})(u_{i}-v)}$
$\leq\frac{4d_{i}}{d(1-u_{i})(u_{i}-v)}+\frac{u_{i}(1-v)\sum_{j=0}^{i-1}x_{j}}{d(1-u_{i})(u_{i}-v)}+\frac{4\sum_{j=0}^{i-1}u_{j}x_{j}}{d(1-u_{i})(u_{i}-v)}$
$\leq\frac{4d_{i-1}}{(1-u_{i})(u_{i}-v)}+\frac{(1-v)\sum_{j=0}^{i-1}x_{j}}{d(1-u_{i})(u_{i}-v)}+\frac{4\sum_{j=0}^{i-1}x_{j}}{d(1-u_{i})(u_{i}-v)}.$
@@
Considering $a_{i}=1+\frac{1}{2^{2^{i}}}$, we have the following:
@@
$\leq 2(u_{i}-v_{i})\sum_{k=0}^{i}2^{g_{k}+\sum_{j=k}^{i}f_{j}}\cdot 4^{i-k+1}$
$\leq 2(u_{i}-v_{i})\cdot(i+1)\cdot 2^{g_{i}}\cdot 2^{\sum_{j=0}^{i}f_{j}}\cdot 4^{i+1}$
$\leq 2(1-v)\cdot(i+1)\cdot d^{2}\cdot 2^{2^{i+1}}\cdot 4^{i+1}$
$\leq 2(1-v)(i+1)\cdot\max\left(d,\frac{1}{1-v}\right)^{6}\cdot 4^{i+1}$
$\leq 2(1-v)((i-1)+3)\cdot\max\left(d,\frac{1}{1-v}\right)^{6}\cdot 4^{i+1}$
$\leq 32\cdot(1-v)\cdot\left(\log\log\max\left(d,\frac{1}{1-v}\right)+3\right)$
$\quad\cdot\max\left(d,\frac{1}{1-v}\right)^{6}\cdot\log^{2}\left(\max\left(d,\frac{1}{1-v}\right)\right).$
@@
$CR$ $\leq\frac{4d_{i-1}}{(1-u_{i})(u_{i}-v)}+\frac{(1-v)\sum_{j=0}^{i-1}x_{j}}{d(1-u_{i})(u_{i}-v)}+\frac{4\sum_{j=0}^{i-1}x_{j}}{d(1-u_{i})(u_{i}-v)}$
$\leq 8\max\left(d,\frac{1}{1-v}\right)^{7}+64\left(\log\log\max\left(d,\frac{1}{1-v}\right)+3\right)$
$\quad\cdot\max\left(d,\frac{1}{1-v}\right)^{12}\cdot\frac{(1-v)^{2}}{d}\cdot\log^{2}\left(\max\left(d,\frac{1}{1-v}\right)\right)$
$+128\left(\log\log\max\left(d,\frac{1}{1-v}\right)+3\right)\cdot\max\left(d,\frac{1}{(1-v)^{2}}\right)^{12}$
$\quad\cdot\frac{1-v}{d}\cdot\log^{2}\left(\max\left(d,\frac{1}{1-v}\right)\right).$
@@ 5.2. Target moving toward the origin
(Theorem 5.3) The optimal competitive ratio follows from having both robots wait at the origin. If we assume that one of the two robots catches up to the target at a distance $d^{\prime}$, then the other robot would be at a distance of $-d^{\prime}$ away from the origin. Alternatively, if both robots are at distance $d^{\prime}$, then the adversary would have placed the target at distance $-d^{\prime}$. In this case, the competitive ratio would be as follows:
@@ 6. Conclusion
Appendix
Appendix A Full Knowledge Model
Proof.
(Theorem 2.1) Robots $R_{1}$ and $R_{2}$ move in opposite directions for a distance of $\frac{d}{1-v}$. If one of the robots, say $R_{1}$, catches the target, then at this point the target would be away from $R_{2}$ by a distance of $\frac{2d}{1-v}$, and $R_{2}$ needs time $\frac{\frac{2d}{1-v}}{1-v}$ to catch the target. Thus, the total time needed to capture the target by both robots is:
$\frac{d}{1-v}+\frac{\frac{2d}{1-v}}{1-v}=\frac{3d-dv}{(1-v)^{2}}.$
This yields the competitive ratio:
$CR=\frac{\frac{3d-dv}{(1-v)^{2}}}{\frac{d}{1-v}}=\frac{3-v}{1-v}.$
This proves Theorem 2.1. ∎
Theorem A.1.
The competitive ratio of any algorithm in the FullKnowledgeAway model is at least $\frac{3-v}{1-v}$.
Proof.
(Theorem A.1) If both robots $R_{1}$ and $R_{2}$ are at the origin and the moving target is at a distance $d$ away from the origin, then any of the two robots needs $\frac{d}{1-v}$ to catch up with the target. Assume $R_{1}$ reaches point $\frac{d}{1-v}$, then the adversary would place the target at $-\frac{d}{1-v}$. At this point, $R_{2}$ catches the target and $R_{1}$ would be far from the target by $\frac{2d}{1-v}$. Thus, for $R_{1}$ to reach the target, it needs time $\frac{\frac{2d}{1-v}}{1-v}=\frac{2d}{(1-v)^{2}}$. The competitive ratio is then:
$CR=\frac{\frac{2d}{(1-v)^{2}}+\frac{d}{1-v}}{\frac{d}{1-v}}=\frac{3-v}{1-v}.$
This proves Theorem A.1. ∎
Proof.
$\frac{d}{v+1}+\frac{2d}{(v+1)^{2}}=\frac{3d+dv}{(v+1)^{2}}.$
The competitive ratio is then:
$CR=\frac{\frac{3d+dv}{(v+1)^{2}}}{\frac{d}{v+1}}=\frac{3+v}{1+v}.$
This proves Theorem 2.3. ∎
Proof.
$CR=\frac{\frac{d}{v}}{\frac{d}{v+1}}=\frac{v+1}{v}.$
This proves Theorem 2.4. ∎
Proof.
  • Case 1: If $v\geq 1$, then clearly $CR\geq 1+\frac{1}{v}$.
This proves Theorem 2.5. ∎
Appendix B NoDistance Model
Proof.
(Theorem 3.1) If the target is at distance $d$ from the origin, then eventually either $R_{1}$ or $R_{2}$ will capture the target. Without loss of generality, assume that $R_{1}$ captures the target. The worst-case scenario occurs if, during iteration $k-1$, $R_{1}$ just misses the target. Specifically, when $R_{1}$ reaches $x_{k-1}$, the target is at distance $x_{k-1}+\epsilon$. Thus, for $R_{1}$ to capture the target, it requires time:
(9) $\frac{x_{k}+vx_{k}}{1-v}.$
Using equations (7), (8), and (9), the competitive ratio is given by:
$=\frac{2x_{k}+d+2\sum_{i=0}^{k-1}x_{i}}{d}$
(10) $=\frac{2\sum_{i=0}^{k-1}x_{i}+2x_{k}}{d}+1.$
To analyze further, assume that in round $k$, the robot captures the target. Then, in round $k-1$, the following must hold:
$\frac{d+\left(2a^{0}+\dots+2a^{k-2}\right)v}{1-v}\geq a^{k-1},$
$\implies d>a^{k-1}-va^{k-1}-2v\frac{a^{k-1}-1}{a-1},$
$\implies d>\frac{a^{k}-a^{k-1}-va^{k}+va^{k-1}-2va^{k-1}+2v}{a-1},$
(11) $\implies d>\frac{a^{k}-a^{k-1}-va^{k}-va^{k-1}+2v}{a-1}.$
Thus, the competitive ratio becomes:
(12) $CR$ $=\frac{2ax_{k}-2}{a^{k}-a^{k-1}-va^{k}-va^{k-1}+2v}+1$ $=\frac{2a^{2}}{a-1-av-v}+1.$
$f^{\prime}(a)$ $=\frac{4a(a-1-av-v)-2a^{2}(1-v)}{(a-1-av-v)^{2}}$
$=\frac{4a^{2}-4a-4va^{2}-4av-2a^{2}(1-v)}{(a-1-av-v)^{2}}$
(13) $=\frac{2a^{2}-2va^{2}-4a-4av}{(a-1-av-v)^{2}}.$
$CR$ $=1+\frac{8(1+v)^{2}}{(1-v)^{2}\cdot(a-1-av-v)}$
$=\frac{8(1+v)^{2}}{(1-v)(1-v^{2})}$
$=\frac{8(1+v)}{1-v^{2}}+1$
(14) $=\frac{(v+3)^{2}}{(1-v)^{2}}.$
This completes the proof of Theorem 3.1. ∎
Proof.
(Theorem 3.2) Let us assume that the target was detected by one of the robots at iteration $k$. Then at the end of iteration $k-1$, both robots would have completed $2(k-1)$ turns. At iteration $k$, one of the robots would make a turn to meet the other robot at the origin, since the other robot already caught the target. Thus, the total number of turns would be $12(k-1)+1=2k-$ 1.
In order to find the number of turns in terms of $d$ and $v$, we assume that the target is caught by one of the robots at iteration $k$. Then, at iteration $k-1$, and based on Equation 11, we have:
$\frac{a^{k-1}(a-1-av-v)}{a}\leq d,$
which implies the following:
$a^{k-1}$ $\leq\frac{ad}{a-1-av-v},$
$\leq\frac{\frac{2d(1+v)}{1-v}}{\frac{2+2v-1+v-2v-2v^{2}-v+v^{2}}{1-v}},$
$=\frac{2d(1+v)}{1-v^{2}},$
$=\frac{2d}{1-v}.$
Taking the logarithm, we obtain:
$k\leq 1+\log\left(\frac{2d}{1-v}\right).$
Thus, the total number of turns would be at most:
$2k-1=1+2\log\left(\frac{2d}{1-v}\right).$
This proves Theorem 3.2. ∎
Proof.
(Proof of Theorem 3.5)
If the target is at distance $d$ away from the origin, then eventually either $R_{1}$ or $R_{2}$ will capture the target. If we assume that $R_{1}$ captures the target, then the worst-case scenario occurs when, during iteration $k-1$, $R_{1}$ just misses the target. In other words, when $R_{1}$ reaches $x_{k-1}$, the target would be at distance $x_{k-1}+\epsilon$. Thus, in order to capture the target, $R_{1}$ needs time:
(17) $\frac{x_{k}-vx_{k}}{1+v}$
Thus, using equations 15, 16, and 17, the competitive ratio is as follows:
$\leq\frac{2\sum_{i=0}^{k-1}x_{i}+2x_{k}+d}{d}$
$=1+\frac{2x_{k+1}-2}{d(a-1)}$
If we assume that in round $k$, the robot captures the target, then in round $k-1$ we should have the following:
$\frac{d-\left(2a^{0}+\dots+2a^{k-2}\right)v}{1+v}\geq a^{k-1}$
$\implies d>a^{k-1}+va^{k-1}+2v\frac{a^{k-1}-1}{a-1}$
Thus, the competitive ratio is as follows:
$CR$ $\leq 1+\frac{2a^{2}\left(\frac{2v+ad-d}{a+av+v-1}\right)-2}{d(a-1)}$
$\leq 1+\frac{4va^{2}+2da^{3}-2da^{2}-2a-2av-2v+2}{d(a-1)(a+av+v-1)}$
$\leq 1+\frac{4av+2da^{2}-2+2v}{d(a+av+v-1)}$
$\leq 1+\lim_{d\to\infty}\frac{4av+2da^{2}-2+2v}{d(a+av+v-1)}$
$=1+\frac{2a^{2}}{a+av+v-1}$
If we set $f(a)=\frac{2a^{2}}{a+av+v-1}$, then
$f^{\prime}(a)$ $=\frac{4a(a+av+v-1)-2a^{2}(1+v)}{(a+av+v-1)^{2}}$
$=\frac{4a^{2}+4va^{2}+4av-4a-2a^{2}-2va^{2}}{(a+av+v-1)^{2}}$
$\leq\frac{2a^{2}+2va^{2}+4av-4a}{(a+av+v-1)^{2}}$
Setting $f^{\prime}(a)=0$ gives $02a+2av+4v-4=$ 0. Thus, $a=\frac{2(1-v)}{1+v}$.
Thus, the competitive ratio would be:
$CR$ $\leq 1+\frac{\frac{8-16v+8v^{2}}{(1+v)^{2}}}{\frac{2-2v^{2}+v+v^{2}-1-v}{1+v}}$
$\leq 1+\frac{8v^{2}-6v+8}{(1+v)(1-v^{2})}$
$=1+\frac{8(1-v)}{(1+v)^{2}}$
This proves Theorem 3.5. ∎
Appendix C No Speed Model
Proof of Theorem 4.3.
In the worst-case scenario, if the target’s speed is very small, the robots may not catch up to the target immediately. In this case, the robots will move a distance $d$ in one direction. At this point, they will be away from the target by a distance of $2d-dv$, where $v$ is the speed of the target.
The competitive ratio (CR) can then be calculated as follows:
$CR$ $=\frac{\text{Distance traveled by the robots}}{\text{Distance traveled by the target}}$
$=\frac{d+\frac{2d-dv}{1+v}}{\frac{d}{1+v}}$
$=3.$
Hence, the competitive ratio is $3$, which proves Theorem 4.3. ∎
Proof of Theorem 4.4.
To prove the lower bound, both points $d$ and $-d$ must be visited by the robots. If only one of these two points, say $-d$, is visited while the other point $d$ is not, the adversary could place the target at $d$. In this case, the robots would fail to catch the target efficiently.
Similarly, if the target is placed at $-d$ and one of the robots visits $d$ first, then by the time the robot reaches $d$, the target would have moved at least $dv$. Following this argument, the competitive ratio would be calculated as follows:
$CR$ $=\frac{\text{Distance traveled by the robots}}{\text{Distance traveled by the target}}$
$=\frac{d+\frac{2d-dv}{1+v}}{\frac{d}{1+v}}$
$=3.$

### v2

@@ 5.1. Target Moving Away from the Origin
$12M^{7}+192(\log\log M+3)\cdot M^{10}\cdot\log^{2}M\cdot\left(\frac{1}{d}\right),$
@@
$\sum_{j=0}^{i-1}x_{j}$ $=\frac{x_{i}(u_{i}-v_{i})-2^{g_{i}}}{v_{i}}$
$\implies\sum_{j=0}^{i}x_{j}=\frac{x_{i+1}(u_{i+1}-v_{i+1})-2^{g_{i+1}}}{v_{i+1}}.$
@@
Thus, based on Equation (18), and considering $f_{i}=2^{2^{i}}$, we have the following:
@@
$CR$ $=\frac{\sum_{j=0}^{i-1}x_{j}+\frac{d_{i}+v\sum_{j=0}^{i-1}x_{j}}{u_{i}-v}+\frac{2\sum_{j=0}^{i-1}u_{i}x_{j}+2u_{i}\left(\frac{d_{i}+v\sum_{j=0}^{i-1}x_{j}}{u_{i}-v}\right)}{1-u_{i}}}{\frac{d}{1-v}}$
$+\frac{\frac{2\sum_{j=0}^{i-1}u_{i}x_{j}+2u_{i}\left(\frac{d_{i}+v\sum_{j=0}^{i-1}x_{j}}{u_{i}-v}\right)+2v\sum_{j=0}^{i-1}u_{i}x_{j}+2vu_{i}\left(\frac{d_{i}+v\sum_{j=0}^{i-1}x_{j}}{u_{i}-v}\right)}{(1-u_{i})(1-v)}}{\frac{d}{1-v}}$
$=\frac{\frac{u_{i}\sum_{j=0}^{i-1}x_{j}+d_{i}}{u_{i}-v}+\frac{2u_{i}\sum_{j=0}^{i-1}u_{i}x_{j}-2v\sum_{j=0}^{i-1}u_{i}x_{j}+2u_{i}d_{i}+2vu_{i}\sum_{j=0}^{i-1}x_{j}}{(1-u_{i})(1-v)(u_{i}-v)}}{\frac{d}{1-v}}$
$+\frac{\frac{2u_{i}\sum_{j=0}^{i-1}u_{i}x_{j}-2v\sum_{j=0}^{i-1}u_{i}x_{j}+2u_{i}d_{i}+2u_{i}v\sum_{j=0}^{i-1}x_{j}+2vu_{i}^{2}\sum_{j=0}^{i-1}x_{j}+2vu_{i}d_{i}}{(1-u_{i})(1-v)(u_{i}-v)}}{\frac{d}{1-v}}$
$\leq\frac{u_{i}\sum_{j=0}^{i-1}x_{j}+d_{i}-u_{i}v\sum_{j=0}^{i-1}x_{j}-vd_{i}-(u_{i})^{2}\sum_{j=0}^{i-1}x_{j}-u_{i}d_{i}+u_{i}^{2}v\sum_{j=0}^{i-1}x_{j}}{d(1-u_{i})(u_{i}-v)}$
$+\frac{d_{i}u_{i}v+2u_{i}\sum_{j=0}^{i-1}u_{i}x_{j}-2v\sum_{j=0}^{i-1}u_{i}x_{j}+2u_{i}d_{i}+2vu_{i}\sum_{j=0}^{i-1}x_{j}}{d(1-u_{i})(u_{i}-v)}$
$+\frac{2u_{i}\sum_{j=0}^{i-1}u_{i}x_{j}-2v\sum_{j=0}^{i-1}u_{i}x_{j}+2u_{i}d_{i}+2u_{i}v\sum_{j=0}^{i-1}x_{j}+2vu_{i}^{2}\sum_{j=0}^{i-1}x_{j}+2vu_{i}d_{i}}{d(1-u_{i})(u_{i}-v)}$
$\leq\frac{u_{i}\sum_{j=0}^{i-1}x_{j}+d_{i}-d_{i}v+3u_{i}d_{i}-u_{i}^{2}\sum_{j=0}^{i-1}x_{j}+3u_{i}^{2}v\sum_{j=0}^{i-1}x_{j}+3d_{i}u_{i}v}{d(1-u_{i})(u_{i}-v)}$
$+\frac{4u_{i}\sum_{j=0}^{i-1}u_{i}x_{j}-4v\sum_{j=0}^{i-1}u_{i}x_{j}-u_{i}v\sum_{j=0}^{i-1}x_{j}+4u_{i}v\sum_{j=0}^{i-1}x_{j}}{d(1-u_{i})(u_{i}-v)}$
$\leq\frac{6d_{i}}{d(1-u_{i})(u_{i}-v)}+\frac{u_{i}(1-v)\sum_{j=0}^{i-1}x_{j}}{d(1-u_{i})(u_{i}-v)}+\frac{6\sum_{j=0}^{i-1}u_{i}x_{j}}{d(1-u_{i})(u_{i}-v)}$
$\leq\frac{6d_{i-1}}{(1-u_{i})(u_{i}-v)}+\frac{(1-v)\sum_{j=0}^{i-1}x_{j}}{d(1-u_{i})(u_{i}-v)}+\frac{6\sum_{j=0}^{i-1}x_{j}}{d(1-u_{i})(u_{i}-v)}.$
@@
Considering $a_{i}=1+\frac{1}{2^{2^{i}}}$ and $g_{i}=2^{i}$, we have the following:
@@
$\leq 4(u_{i}-v_{i})\sum_{k=0}^{i}2^{g_{k}+\sum_{j=k}^{i}f_{j}}\cdot 4^{i-k}$
$\leq 4(u_{i}-v_{i})\cdot(i+1)\cdot 2^{g_{i}}\cdot 2^{\sum_{j=0}^{i}f_{j}}\cdot 4^{i}$
$\leq 4\frac{v_{i}}{2^{2^{i}}}\cdot(i+1)\cdot d^{2}\cdot 2^{2^{i+1}}\cdot 4^{i}$
$\leq 4(i+1)\cdot\max\left(d,\frac{1}{1-v}\right)^{4}\cdot 4^{i}$
$\leq 4((i-1)+3)\cdot\max\left(d,\frac{1}{1-v}\right)^{4}\cdot 4^{i}$
$\leq 16\cdot\left(\log\log\max\left(d,\frac{1}{1-v}\right)+3\right)$
$\quad\cdot\max\left(d,\frac{1}{1-v}\right)^{4}\cdot\log^{2}\left(\max\left(d,\frac{1}{1-v}\right)\right).$
@@
$CR$ $\leq\frac{6d_{i-1}}{(1-u_{i})(u_{i}-v)}+\frac{(1-v)\sum_{j=0}^{i-1}x_{j}}{d(1-u_{i})(u_{i}-v)}+\frac{6\sum_{j=0}^{i-1}x_{j}}{d(1-u_{i})(u_{i}-v)}$
$\leq 12\max\left(d,\frac{1}{1-v}\right)^{7}+32\left(\log\log\max\left(d,\frac{1}{1-v}\right)+3\right)$
$\quad\cdot\max\left(d,\frac{1}{1-v}\right)^{10}\cdot\frac{(1-v)}{d}\cdot\log^{2}\left(\max\left(d,\frac{1}{1-v}\right)\right)$
$+192\left(\log\log\max\left(d,\frac{1}{1-v}\right)+3\right)\cdot\max\left(d,\frac{1}{(1-v)^{2}}\right)^{10}$
$\quad\cdot\frac{1}{d}\cdot\log^{2}\left(\max\left(d,\frac{1}{1-v}\right)\right).$
@@ 5.2. Target moving toward the origin
(Theorem 5.3) The optimal competitive ratio follows from having both robots wait at the origin. If we assume that one of the two robots catches up to the target at a distance $d^{\prime}$, then the other robot would be at a distance of $d^{\prime}$ away from the origin. Alternatively, if both robots are at distance $d^{\prime}$, then the adversary would have placed the target at distance $d^{\prime}$. In this case, the competitive ratio would be as follows: